Virtual memory address translation: a) Consider a machine with a physical memory of 8 GB, a page size of 4 KB, and a page table entry size of 4 bytes. How many levels of page tables would be required to map a 52-bit virtual address space if every page table fits into a single page? b) Without a cache or TLB, how many memory operations are required to read or write a page in physical memory? c) How much physical memory is needed for a process with three pages of virtual memory?
Answer #1

Answer a

Physical memory size = 8 GB = 233 Bytes

Page Size = 4 KB = 212 Bytes

Page Table Entry Size = 4 Bytes => #Entries in page table = 212/4 = 210

Logical Address Space = 252 Bytes => #Pages = 252/212 = 240 => #Entries = 240

# First Level Page tables = Total entries / #entries in one page table = 240/210 = 230

# Second Level Page tables = Total entries of first level / #entries in one page table = 230/210 = 220

# Third Level Page tables = Total entries of second level / #entries in one page table = 220/210 = 210​​​​​​​

# Fourth Level Page table = Total entries of third level / #entries in one page table = 210/210 = 1​​​​​​​

So total 4 level of page tables required.

Answer b

Without a cache or TLB, There will be 4 memory accesses to read page tables to get frame number and then 1 memory access for actual data. => Total 5

Answer c

Physical memory = 3*4 = 12 KB